OpenStudy (anonymous):

Solve for j. -9j = -8j + 8

OpenStudy (mollynicole2015):

Simplifying -9j = -8j + 8 Reorder the terms: -9j = 8 + -8j Solving -9j = 8 + -8j Solving for variable 'j'. Move all terms containing j to the left, all other terms to the right. Add '8j' to each side of the equation. -9j + 8j = 8 + -8j + 8j Combine like terms: -9j + 8j = -1j -1j = 8 + -8j + 8j Combine like terms: -8j + 8j = 0 -1j = 8 + 0 -1j = 8 Divide each side by '-1'. j = -8 Simplifying j = -8 So how do we know this is true?

OpenStudy (anonymous):

Put the answer back into the equation?

OpenStudy (mollynicole2015):

Right! So do that. Replace j with -8

OpenStudy (misty1212):

lol HI!!

OpenStudy (misty1212):

where in the world did that come from?

OpenStudy (anonymous):

-9 x -8 = -8 x -8 + 8 -9 x -8 = 72 -8 x -8 + 8 = 72 72 = 72 So our equation is true.

OpenStudy (anonymous):

Hi.

OpenStudy (mollynicole2015):

Hey.

OpenStudy (misty1212):

\[-9j=-8x+8\] add \(8j\) get \[-j=8\] change the sign get \[j=-8\] that is all

OpenStudy (mollynicole2015):

Great job, @ArianatorNo1

OpenStudy (anonymous):

Thanks all!